The reaction between H2 and-Br2 to form HBr in the presence of light is initiated by the photon decomposition of Br2 into free bromine atoms (free radicals) by absorption of light The bond dissociation energy of Br2 is 192 kJ/mole. What is the longest wavelength of the photon that would initiate the reaction?

Grade:

2 Answers

siddharth gupta
9 years ago
BOND DISSOCIATION ENTHALPY IS THE MINIMUM AMOUNT OF ENERGY REQUIRED TO INITIATE BOND DISSOCIATION.
E=hc/l where l is the wavelength of light required.
for constant hc; l is inversely proportional to E
i.e maximum wavelength for minimum energy=hc/E=((6.626x3x10-26)x6.02 x 1023)/192 x 103 J=6.2 X 10-7m
 
needhelp
9 years ago
Use de broglie wavelength eqn
E=hc/lamda